OpenStudy (anonymous):

Help Please! Metal. Write the quadratic function in vertex form. y = x2 + 4x + 7 y = (x - 2)2 + 3 y = (x - 2)2 - 3 y = (x + 2)2 - 3 y =(x + 2)2 + 3

OpenStudy (anonymous):

There are four options.

OpenStudy (jskhupmang):

ok do know how to write general form?

OpenStudy (jskhupmang):

@clairvoyant1

OpenStudy (anonymous):

No.

OpenStudy (jskhupmang):

which one do you what me to explain?

OpenStudy (anonymous):

yes, please.

OpenStudy (jskhupmang):

first I would use perfect sqaure property y= x^2+4x+4+7-4

OpenStudy (jskhupmang):

eventhough I added 4 I see that I subtract 4 right? So adding 4 really means nothing

OpenStudy (jskhupmang):

I meant You*

OpenStudy (jskhupmang):

y=(x-2)^2+3 so now you can factor so the vertex is 2,3

OpenStudy (jskhupmang):

Did you know how I got the vertex? @clairvoyant1

OpenStudy (jskhupmang):

the factor (x-2) is nothing more than x=2. It is always the opposite of whats inside hope this help

OpenStudy (jskhupmang):

although the 3 outside of the function is not affected so the sign does not change

OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is h = -16t2 + 672t. After how many seconds does it reach its maximum height? 21 s 31.5 s 42 s 10 s

OpenStudy (jskhupmang):

I think the fuction you gave me is wrong I meant there should be a constant right? after 672t

OpenStudy (jskhupmang):

in my graphing calculator it does not fall down. It goes all the way up

OpenStudy (anonymous):

No, that's the problem.

OpenStudy (jskhupmang):

so I believe the function is -16t^2+672t+t? when you say t seconds

OpenStudy (anonymous):

I believe so.

OpenStudy (jskhupmang):

ok give me a sec

OpenStudy (anonymous):

Ok.

OpenStudy (jskhupmang):

aw now i see the greatest seconds would have the greatest maximum! right!? lol is that easy

OpenStudy (jskhupmang):

since its never fall down

OpenStudy (anonymous):

21 s?

OpenStudy (jskhupmang):

sorry i was wrong Now I got it

OpenStudy (anonymous):

That's Ok, I'm really concerned about this problem.

OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote of y = 2/9.

OpenStudy (jskhupmang):

yes you were right 21s

OpenStudy (jskhupmang):

You dope me! lol

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

That's like an Essay question.

OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote of y = 2/9.

OpenStudy (jskhupmang):

hmm so they were asking you a rational function

OpenStudy (anonymous):

yea

OpenStudy (jskhupmang):

ok this is an example of y= 2/3. the function is f(x)=(2x-11)/(3x+5)

OpenStudy (jskhupmang):

basically the it's the coefficient of the function

OpenStudy (jskhupmang):

Can you see that 2xWC divide by 3xWC WC means whocare

OpenStudy (anonymous):

Lol, I see that clearly.

OpenStudy (jskhupmang):

Yup :)

OpenStudy (anonymous):

State the domain of the rational function. f(x) = seven divided by quantity fourteen minus x.

OpenStudy (anonymous):

State the domain of the rational function. f(x) = seven divided by quantity fourteen minus x. A) All real numbers except 14. B) All real numbers except -14 and 14. C) All real numbers except 7. D) All real numbers except -7 and 7.

OpenStudy (jskhupmang):

aw so we have to solve and what ever that is can not be equal to x.

OpenStudy (anonymous):

\[f(x) = \frac{ 7 }{ 14-x}\]

OpenStudy (jskhupmang):

so let me try to solve it

OpenStudy (jskhupmang):

x can not be equal to 14

OpenStudy (jskhupmang):

hold on though i wanna make sure we dont have to simplify 7 and 14

OpenStudy (anonymous):

So all real number except 14?

OpenStudy (jskhupmang):

yes

OpenStudy (jskhupmang):

when they say state the domain, they are just really saying solve the dominator of the function

OpenStudy (anonymous):

State how many imaginary and real zeros the function has. f(x) = x5 + 7x4 + 2x3 + 14x2 + x + 7 3 imaginary; 2 real 4 imaginary; 1 real 0 imaginary; 5 real 2 imaginary; 3 real

OpenStudy (jskhupmang):

c

OpenStudy (jskhupmang):

does not have imaginary because theres no missing order of power ya know what im sayin? lol

OpenStudy (jskhupmang):

like the 5 4 3 2 1

OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 3(x + 8)2(x - 8)3 -8, multiplicity 2; 8, multiplicity 3 4, multiplicity 1; 8, multiplicity 1; -8, multiplicity 1 -8, multiplicity 3; 8, multiplicity 2 4, multiplicity 1; -8, multiplicity 3; 8, multiplicity 3

OpenStudy (jskhupmang):

I had a hard time looking at the function was it the second power lol that confused me

OpenStudy (anonymous):

That's Ok.

OpenStudy (jskhupmang):

no i know it now

OpenStudy (jskhupmang):

its a

OpenStudy (jskhupmang):

the powers are count as multiplicity and the factors were zeros

OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all potential rational zeros f(x) = x3 - 10x2 + 4x - 24 +/-1, +/-2, +/-3, +/-4, +/-6, +/-8, +/-12, +/-24 +/-1, +/-One divided by two., +/-2, +/-3, +/-4, +/-6, +/-8, +/-12, +/-24 +/-1, +/-2, +/-3, +/-4, +/-24 +/-1, +/-2, +/-3, +/-4, +/-6, +/-12, +/-24

OpenStudy (jskhupmang):

are they the answers? like abcde

OpenStudy (anonymous):

yea

OpenStudy (jskhupmang):

whatever is a multiplicity of 24 so

OpenStudy (jskhupmang):

which one do you think Brittany? lol Thanks for being my fan love ya

OpenStudy (anonymous):

"a" your awesome, where are u from?

OpenStudy (jskhupmang):

Im from Florida How about you?

OpenStudy (anonymous):

Florida!

OpenStudy (jskhupmang):

wow! great! hope to meet someday lol

OpenStudy (jskhupmang):

what school do you go to?

OpenStudy (anonymous):

I'm in the orlando area, so are you in college?

OpenStudy (jskhupmang):

nope Im still high school

OpenStudy (jskhupmang):

senior I go to Monarch High have you ever heard of it?

OpenStudy (anonymous):

Lol, me too. I go to Lake Brantley... every heard of it?

OpenStudy (jskhupmang):

no lol

OpenStudy (anonymous):

I haven't heard of Monarch I'll look it up right now. One Sec.

OpenStudy (jskhupmang):

I came in fl when I was 14 so I have never heard of it

OpenStudy (jskhupmang):

sure take your time

OpenStudy (anonymous):

Oh ok. you should look up Lake Brantley, Check us Out.

OpenStudy (anonymous):

I'm a Senior to Btw.

OpenStudy (jskhupmang):

Yea i saw it. Its in orlando wow thats very far from me

OpenStudy (anonymous):

Lol, yeah

OpenStudy (jskhupmang):

Yay! haha

OpenStudy (jskhupmang):

So are taking online classes?

OpenStudy (jskhupmang):

you*

OpenStudy (anonymous):

Yeah, a few.

OpenStudy (anonymous):

so let's get this assignment done then we'll talk.

OpenStudy (jskhupmang):

ok Im ready for you anytime :)

OpenStudy (anonymous):

Lol ok, One Sec.

OpenStudy (anonymous):

\[\frac{ 2x ^{3 }- 3x ^{2}- 5x + 4 }{ x - 2 }\]

OpenStudy (anonymous):

Divide using synthetic division, and write a summary statement in fraction form.

OpenStudy (anonymous):

I'm typing up the answer choices.

OpenStudy (jskhupmang):